# Printing and Scanning Attack for Image Counter Forensics

### v1

@@ Abstract
Examining the authenticity of images has become increasingly important as manipulation tools become more accessible and advanced. Recent work has shown that while CNN-based image manipulation detectors can successfully identify manipulations, they are also vulnerable to adversarial attacks, ranging from simple double JPEG compression to advanced pixel-based perturbation. Rephrase(hailey): In this paper we explore another method of highly plausible attack: printing and scanning. We demonstrate the vulnerability of two state-of-the-art models to this type of attack. We also propose a new machine learning model that performs comparably to these state-of-the-art models when trained and validated on printed and scanned images. Of the three models, our proposed model performs the best when trained and validated on images from a single printer. To facilitate this exploration, we create a dataset of over 6,000 printed and scanned image blocks. Further analysis suggests that variation between images produced from different printers is significant, large enough that good validation accuracy on images from one printer does not imply similar validation accuracy on images from a different printer.
@@ 1 Introduction
Determining the authenticity of an image is becoming increasingly important for legal proceedings, criminal investigations, and verifying identity-supporting documents. In recent years, Convolutional Neural Networks have been employed to detect image manipulations, ranging from identifying splicing and copy-move forgeries [30] to manipulations such as contrast enhancement [32, 38], resampling [28], JPEG compression [14], Gaussian blurring [19, 4], median filtering [5], and additive white Gaussian noise [2]. Of this latter group, some may be innocuously applied but often are maliciously included in post-processing to disguise other manipulations [6], .
@@ 2 Related Work
As the scope of this paper is limited to CNN-based global image manipulation detectors, we provide background on convolutional neural networks and CNN-based detectors. We similarly provide context on adversarial attacks on CNNs generally as well as specifically on CNN-based image manipulation detectors. Related to this work is work on detecting manipulations through inconsistencies in lighting [20] and despite various compression qualities [13]. [20] contributes significantly to this problem area, though without examining models that leverage deep learning. [13] similarly advances the area, but without addressing specifically the problem of printing and scanning in relation to CNN-based detectors, and is thus complementary to this work.
2.1 Convolutional Neural Networks
Convolutional neural networks, and deep learning approaches more generally, are an extension of neural networks. The architecture is defined by a set of parameters used to design the network and generally consists of stacked layers. This approach has proven effective in extracting higher-order features, particularly in object identification, but recent work has also demonstrated that these types of networks, with some modifications, are also useful for extracting the low level features useful for detecting global image manipulations. In a Convolutional Neural Network (CNN), the first layer is a set of convolutional feature extractors that convolve across the image in parallel using a set of learnable filters. The filters are initialized randomly, and the values are updated iteratively through back-propagation [22, 23]. The filters use a sliding window approach to reach all regions of the input image, producing feature maps. The following hidden layers function similarly, convolving over the resulting feature maps to produce higher-level feature maps. Each convolutional layer is followed by an activation function to provide non-linearity. Additionally, convolutional layers are often followed by pooling layers (max, average, stochastic, etc) to reduce the dimensionality of the resulting features and reduce over-fitting. The output of the final convolutional layer is inputted into a fully-connected neural network for classification.
@@ 2.2 Deep Learning for Image Forensics
Table 1: Descriptions and sizes of each dataset used for training and validation. Size refers to the number of 256x256 or 299x299 image blocks in each dataset, which is then split in 75% training and 25% validation. X1 and X2 refer to Xerox1 and Xerox2 printers (Section 4.1), respectively. The labels refer to the labels used when training and evaluating on each dataset
@@ 2.4 Adversarial Attacks in Image Forensics
While CNN-based classifiers have achieved high performance on benchmark image forensic tasks, recent research in computer vision has demonstrated that CNN-based detectors, like CNNs more broadly, are highly vulnerable to adversarial attacks. For example, in [21], the authors demonstrate that a GAN-based architecture can conceal 3x3 median filtering manipulation, causing a detector to label the image as non-manipulated, including for the CNN-based detectors proposed in [2] and [9]. Additionally, a method of adversarial attack based on small pixel-based distortions has been proposed for fooling global image manipulation detectors [35]. However, [36] notes that unlike in most pattern recognition tasks, pixel-based adversarial attacks such as Fast Gradient Sign Method (FGSM) [17] and Jacobian-based Saliency Map Attack (JSMA) [27], are not for the most part transferable between models.
@@ 3 Models
Here we describe our proposed model architecture for improved performance on printed and scanned images. We compare our model to the model proposed in [2], the inspiration for the constrained convolutional layer, as well as to XceptionNet (Xception) [11], as they have nearly identical number of parameters and similar architecture, so the difference in performance cannot be attributed to increased network capacity.
@@ 3.2 Bayar2016
Proposed in 2016, the constrained convolution method of image manipulation detection, hereafter referred to at Bayar2016, proposes a three-layer CNN, with two max-pooling layers and three fully-connected layers (including the initial constrained convolutional layer) [2]. This model demonstrated impressive results in discerning between the six manipulations investigated in this paper using the dataset described in the next section, achieving 99.9% validation accuracy.
@@ 3.3 Xception
In addition to a the Bayar2016 shallow network, recent work has demonstrated that increasing network depth can dramatically improve model generalization. To compare with a model of similar depth that also uses separable convolutional layers, we experiment with XceptionNet, a deep network comprising of 42 layers, including separable convolutional layers [11]. The network design is built upon Inception architecture [33], with the innovation of separable filters.
@@ 4 Datasets
Following the procedure described in [2], we use images from the first IEEE IFSTC Image Forensics Challenge [29], obtaining 3,334 images of size 1024x768, which was further split into training, validation and testing data. The dataset comprises of many images captured from several different digital cameras of both indoor and outdoor scenes.
@@ 4.1 Printing and Scanning
We used three different printers and one scanner to create a dataset of printed and scanned images: one Dell S3845CDN Laser Multifunction Printer, one Xerox Altalink C8070 Multifunction Printer, and one Xerox WorkCentre 7970 Multifunction Printer, which we refer to as Dell, Xerox1 and Xerox2 respectively hereafter. We printed 50 images of each manipulation type on each printer and used the Dell scanner to scan each image (see Figure 1). After scanning and extracting the images from the resulting pdfs, the image sizes were 1700 x 2200 pixels, which was then center-cropped to 1536 x 1792 to remove the white border added by the scanning process. Each image was then split into 42 256x256 blocks (or 299x299 blocks for our proposed model and for Xception), resulting in 2,142 image blocks of each class from each printer (see Figure 2). We limited our data creation to 900 full-page color images both for budget constraints and environmental concerns, thus presenting synthetic printing and scanning simulation as an avenue of possible future work.
@@ 4.2 Manipulations
We post-processed each image with additive white Gaussian noise (AWGN), Gaussian blurring (GB), JPEG compression (JPEG), median filtering (MF) and re-sampling (RS) as well as retaining the pristine image (PR), yielding six classes, per the procedure described in [2].
Additive white Gaussian noise constructs a noise matrix of the same shape as the image according to a normal distribution with a given sigma value and adds this matrix to the original image. The result is then normalized to values between 0 and 255. Gaussian blurring blurs the image using a Gaussian filter by convolving the input image using a given kernal. JPEG compression is a lossy compression method which compresses the image through converting the color map, down-sampling and Discrete Cosine Transform (DCT). Median filtering replaces each pixel with the median value of the neighboring pixels using a given kernal area. Bilinear resampling works similarly, resizing the image using the distance-weighted average of the neighboring pixels to estimate the new pixel value. See Table 2 for parameter details.
@@ 5 Experiments
We trained each model (our proposed model, Bayar2016, and Xception) on a variety of training sets and evaluated each trained model on a variety of validation datasets (See Table 1).
We first evaluated the extent to which our selected models can correctly classify the validation images after printing and scanning. We trained each model on the original dataset (before printing and scanning) with all six classes: additive white gaussian noise (AWGN), gaussian blurring (GB), JPEG compression (JPEG), median filtering (MF), bilinear resampling (RS) and pristine or no manipulation (PR). For a more complete analysis, we removed the bilinear resampling (RS) and JPEG compression (JPEG) classes from the training and validations sets and retrained the models, as these two classes could intuitively be considered similar to changes introduced during the printing and scanning process (see Table 3).
Second, we explored countering this vulnerability by training on the printed and scanned image blocks [17]. We trained each model on the printed and scanned image blocks from a single printer. The dataset (see Table 1, Xerox1) consists of 50 full images (1700 x 2200 pixels), which were then divided into 256x256 for the Bayar2016 model or 299x299 for the Xception models. This resulted in 2,142 image blocks for each dataset, which was divided into training and validation sets of size 1722 and 420 respectively, using only the central images to avoid including border artifacts from the scanning process.
@@ 7 Summary
We investigated the robustness of current state-of-the-art image manipulation detection models in the context of printing and scanning, and found that these state-of-the-art models perform poorly on printed and scanned image data. We proposed our own novel model architecture, which performs  5% better than the state-of-the-art models when trained and evaluated on images from a single printer. We constructed a dataset of over 6,000 printed and scanned image blocks which we plan to release to the community for further investigation.

### v2

@@ Abstract
Examining the authenticity of images has become increasingly important as manipulation tools become more accessible and advanced. Recent work has shown that while CNN-based image manipulation detectors can successfully identify manipulations, they are also vulnerable to adversarial attacks, ranging from simple double JPEG compression to advanced pixel-based perturbation. In this paper we explore another method of highly plausible attack: printing and scanning. We demonstrate the vulnerability of two state-of-the-art models to this type of attack. We also propose a new machine learning model that performs comparably to these state-of-the-art models when trained and validated on printed and scanned images. Of the three models, our proposed model outperforms the others when trained and validated on images from a single printer. To facilitate this exploration, we create a dataset of over 6,000 printed and scanned image blocks. Further analysis suggests that variation between images produced from different printers is significant, large enough that good validation accuracy on images from one printer does not imply similar validation accuracy on identical images from a different printer.
@@ 1 Introduction
Determining the authenticity of an image is becoming increasingly important for legal proceedings, criminal investigations, and verifying identity-supporting documents. In recent years, Convolutional Neural Networks (CNNs) have been employed to detect image manipulations, ranging from identifying splicing and copy-move forgeries [28] to manipulations such as contrast enhancement [30, 36], resampling [26], JPEG compression [14], Gaussian blurring [19, 4], median filtering [5], and additive white Gaussian noise [2]. Of this latter group, some may be either innocuously applied or maliciously included [6]. [^1] Global manipulations such as these can be used on their own to obscure features of an image or used in tandem with another type of manipulation such as copy-move or splicing. We adopt the approach used other researchers in this area in investigating these types of global image manipulations alone (See Section 2).
@@ 2 Related Work
As this paper primarily investigates manipulation detectors based on convolutional neural networks (CNNs), we provide background on CNN-based manipulation detectors. Similarly, we provide context on adversarial attacks on CNNs generally as well as specifically on CNN-based image manipulation detectors.
Related to this work is work on detecting manipulations through inconsistencies in lighting [20] and despite various compression qualities [13]. Additionally, [20] contributes significantly to this problem area, though without examining models that leverage deep learning. [13] explores a similar problem, but without addressing specifically the problem of printing and scanning in relation to CNN-based detectors, and is thus complementary to this work.
@@ 2.1 Deep Learning for Image Forensics
The performance achieved by constrained convolutional layers and particularly deep networks is particularly impressive. These techniques serve as inspiration for our proposed model, and we thus compare our proposed model with models that leverage these modifications.
Table 1: Descriptions and sizes of each dataset used for training and validation. Size refers to the number of 299x299 or 256x256 image blocks in each dataset, which is then split in 75% training and 25% validation. X1 and X2 refer to Xerox1 and Xerox2 printers (Section 4.1), respectively. The labels refer to the labels used when training and evaluating on each dataset
@@ 2.3 Adversarial Attacks in Image Forensics
While CNN-based classifiers have achieved high performance on benchmark image forensic tasks, recent research in computer vision has demonstrated that CNN-based manipulation detectors, like CNNs more broadly, are highly vulnerable to adversarial attacks. For example, in [21], the authors demonstrate that a GAN-based architecture can conceal 3x3 median filtering manipulation, one of the manipulations we explore in this paper. This type of adversarial attack causes a detector to label the image as non-manipulated, including for the CNN-based detectors proposed in [2] and [9]. Additionally, a method of adversarial attack based on small pixel-based distortions has been proposed for fooling global image manipulation detectors [33]. However, [34] notes that unlike in most pattern recognition tasks, pixel-based adversarial attacks such as Fast Gradient Sign Method (FGSM) [17] and Jacobian-based Saliency Map Attack (JSMA) [25], are not for the most part transferable between manipulation detection models.
@@ 3 Models
Here we describe our proposed model architecture for improved performance on printed and scanned images. We compare our model’s performance with the model proposed in [2], the inspiration for the constrained convolutional layer. We additionally compare our model with XceptionNet (Xception) [11], as it and our proposed model have nearly identical number of parameters and similar architecture, so the difference in performance cannot be attributed to increased network capacity.
@@ 3.2 Bayar2016
Proposed in 2016, the constrained convolution method of image manipulation detection, hereafter referred to at Bayar2016, proposes a three-layer CNN, with two max-pooling layers and three fully-connected layers (including the initial constrained convolutional layer) [2]. This model demonstrates impressive results in discerning between the six manipulations investigated in this paper using the dataset described in the next section, achieving 99.9% validation accuracy.
@@ 3.3 Xception
In addition to a the Bayar2016 shallow network, recent work has demonstrated that increasing network depth can dramatically improve model generalization. To compare with a model of similar depth that also uses separable convolutional layers, we experiment with XceptionNet, a deep network comprising of 42 layers, including separable convolutional layers [11]. The network design is built upon Inception architecture [31], with the innovation of separable filters. Similar to Bayar2016, this model also achieves near 99% accuracy on the dataset described in [2] before printing and scanning.
@@ 4 Datasets
For accurate comparison, we follow the procedure described in [2], using images from the first IEEE IFSTC Image Forensics Challenge [27]. The portion of the dataset used consists of 3,334 images of size 1024x768, which was further split into training, validation and testing data. The images are captured from several different digital cameras of both indoor and outdoor scenes.
@@ 4.1 Printing and Scanning
We used three different printers and one scanner to create a dataset of printed and scanned images: one Dell S3845CDN Laser Multifunction Printer, one Xerox Altalink C8070 Multifunction Printer, and one Xerox WorkCentre 7970 Multifunction Printer, which we refer to as Dell, Xerox1 and Xerox2 respectively hereafter. We printed 50 images of each manipulation type on each printer and used the Dell scanner to scan each image (see Figure 1). After scanning and extracting the images from the resulting pdfs, the image sizes were 1700 x 2200 pixels, which was then center-cropped to 1536 x 1792 to remove the white border added by the scanning process. Each image was then split into 42 299x299 blocks (or 256x256 blocks for Bayar2016), resulting in 2,142 image blocks of each class from each printer (see Figure 2). We limited our data creation to 900 full-page color images both for budget constraints and environmental concerns; creating a synthetic dataset through printing and scanning simulation may be an avenue of future work.
@@ 4.2 Manipulations
Again following the procedure described in [2], we manipulated each image with each of six manipulation types: additive white Gaussian noise (AWGN), Gaussian blurring (GB), JPEG compression (JPEG), median filtering (MF), re-sampling (RS) and retaining the pristine image (PR).
  • Additive white Gaussian noise constructs a noise matrix of the same shape as the image according to a normal distribution with a given sigma value and adds this matrix to the original image. The result is then normalized to values between 0 and 255.
  • Gaussian blurring blurs the image using a Gaussian filter by convolving the input image using a given kernal.
  • JPEG compression is a lossy compression method which compresses the image through converting the color map, down-sampling and Discrete Cosine Transform (DCT).
  • Median filtering replaces each pixel with the median value of the neighboring pixels using a given kernal area.
  • Bilinear resampling works similarly, resizing the image using the distance-weighted average of the neighboring pixels to estimate the new pixel value.
See Table 2 for manipulation parameter details.
@@ 5 Experiments
We trained each model (our proposed model, Bayar2016, and Xception) on a variety of training sets and evaluated each trained model on multiple validation datasets (See Table 1).
We first investigated the extent to which our selected models can correctly classify the validation images after printing and scanning. We trained each model on the original dataset (before printing and scanning) with all six classes: additive white gaussian noise (AWGN), gaussian blurring (GB), JPEG compression (JPEG), median filtering (MF), bilinear resampling (RS) and pristine or no manipulation (PR). For a more complete analysis, we removed the bilinear resampling (RS) and JPEG compression (JPEG) classes from the training and validations sets and retrained the models, as these two classes could intuitively be considered similar to changes introduced during the printing and scanning process (see Table 3).
Second, we explored countering this vulnerability by training on the printed and scanned image blocks [17]. We trained each model on the printed and scanned image blocks from a single printer. The dataset (see Table 1, Xerox1) consists of 50 full images (1700 x 2200 pixels), which were then divided into 299x299 for our proposed model and Xception, and 256x256 for the Bayar2016 model. This resulted in 2,142 image blocks for each dataset, which was divided into training and validation sets of size 1722 and 420 respectively, using only the central images to avoid including border artifacts from the scanning process.
@@ 7 Summary
We investigated the robustness of current state-of-the-art image manipulation detection models in the context of printing and scanning, and found that these models perform poorly on printed and scanned image data. We proposed our own novel model architecture, which performs  5% better than the state-of-the-art models when trained and evaluated on images from a single printer. We constructed a dataset of over 6,000 printed and scanned image blocks which we plan to release to the community for further investigation.